# Sylber: Syllabic Embedding Representation of Speech from Raw Audio

### v1

@@ 7 Conclusion
We propose a novel self-supervised learning framework of speech that learns to transform speech waveform to syllabic embedding. As results, we achieve a state-of-the-art unsupervised syllable detection and discovery performance with a significant reduction in time complexity of segmentation algorithm, down to $O(n)$. The resynthesis experiments validate that our model can provide syllabic tokens that encodes intelligible speech content in highly efficient way. Up to our knowledge this is the first demonstration of speech synthesis from syllable-level tokens. Furthermore, a linguistic-theory inspired analysis reveals that Sylber also segments the embedding space, showing emergent categorical perception of speech. To sum, we introduce a new method for representing speech features as syllables, offering promising potential for more efficient speech tokenization and spoken language modeling.
Limitations Our model is not yet suitable for universal speech representation, which the most speech SSL approaches aim for (Yang et al., 2021). We find that Sylber degrades in some SUPERB downstream tasks, which we believe, is due to the parsimonious structure we are imposing (Appendix A.3). Furthermore, our model has not been scaled to a larger speech corpus, which may provide more gains than our current setting, as shown by previous studies (Borsos et al., 2022; Hassid et al., 2024). We leave these for future directions to explore.
Acknowledgments
This research is supported by the following grants to PI Anumanchipalli — NSF award 2106928, BAIR Commons-Meta AI Research, the Rose Hills Innovator Program, UC Noyce Initiative at UC Berkeley, and Google Research Scholar Award. Special thanks to Shang-Wen (Daniel) Li and Abdelrahman Mohamed for valuable discussions and advice.
@@ A.1.1 Self-segmentation distillation
Given a speech audio, $x$, we extract features, $M_{S}(x)=z^{S}$ and $M_{T}(x)=z^{T}$, where $M_{S}$ and $M_{T}$ are the student and teacher models, respectively. The unsupervised segmentation algorithm, Useg, outputs segment boundaries from $z$ as $\text{Useg}(z)=\{s\}^{N}$, where $N$ is the number of discovered segments, and $s\in\mathbb{N}^{2}$ denotes start and end frames of the segment, indexed as $s_{j,0}$ and $s_{j,1}$ for the $j$-th segment. We define an assignment function, $A(i)=j$, that gives the index of the segment, $j$, given a frame number, $i$, such that $s_{j,0}\leq i<s_{j,1}$. When there is no assignable segment, $A(i)=-1$, meaning $i$ is a non-speech frame. The segment-averaged feature, $v_{j}$, is defined by averaging across frames in the $j$-th segment, $v_{j}=\frac{1}{p-q}\sum_{k\in[p,q]}z_{k}$, where $(p,q)=(s_{j,0},s_{j,1})$. Then, $v^{T}_{A(i)}$ indicates the teacher’s segment-averaged feature of the segment that $i$-th frame belongs to, being the target of the regression, which is zero for non-speech frame, $v^{T}_{-1}=0$. Finally, the loss function of the proposed self-segmentation distillation is defined as $\mathcal{L}_{\text{SegDistill}}:=\sum_{i}||v^{T}_{A(i)}-z^{S}_{i}||_{2}^{2}$.
@@
5:         if $\text{speech}_{i}$ and (no current segment or $\text{sim}(s_{i},s_{i-1})<M_{thr}$then
@@
8:              Add $s_{i}$ to current segment $S_{k}$
@@
15:              Define local search range from $a=\text{midpoint}(S_{k})$ to $b=\text{midpoint}(S_{k+1})$
16:              Find optimal boundary $j^{*}$ = $\operatorname*{arg\,max}_{j}\sum_{i=a}^{j}\text{sim}(s_{i},\text{avg}(S_{k}))+\sum_{i=j+1}^{b}\text{sim}(s_{i},\text{avg}(S_{k+1}))$
17:              Update boundary to $j^{*}$
@@
19:     end for
20:     return $S$
21:end procedure
A.1.2 Noise Augmentation
@@
For denoising objective, we mix the input with a randomly sampled environmental sound or other speech audio. For mixing with environmental sound, we randomly select a clip from Reddy et al. (2021) and sample a 5 seconds clip from it. We first z-score the waveform and multiply by a factor sampled from $[0.05,0.7]$, and mix with original speech audio. Note that the original speech is also z-scored. For mixing with other speech, we randomly select another clip in the batch and shift from left or right with a percentage sampled from $[0.4,0.7]$, to make sure the original speech holds the dominant information context in the mixture. The magnitude is also modulated by multiplying by a factor sampled from $[0.0,0.2]$. We apply this augmentation to 20% of the samples in the batch, and only to the inputs fed to the student model. Within the 20%, we have the source of noise be 75% environmental noise and 25% other speech.
A.1.3 Token-to-speech
@@
Articulatory Encodec Articulatory Encodec (Cho et al., 2024c) is composed of articulatory encoding and decoding. The encoding pipeline outputs 14 articulatory features at 50 Hz are used, which are composed of the XY coordinates of 6 articulators (lower incisor; upper and lower lips; tongue tip, blade and dorsum;), and loudness and pitch. These are interpretable and grounded representations of speech that are fully informative of speech contents (Cho et al., 2024c). The decoder, or articulatory vocoder, is a Hifi-GAN (Kong et al., 2020) conditioned on a speaker embedding inferred from a separate speaker encoder. Cho et al. (2024c) shows that Articulatory Encodec successfully decomposes speech contents and speaker identity, by normalizing pitch to remove speaker specific pitch level. We replicate the implementation from Cho et al. (2024c), except that we change the layer of WavLM from which speaker information is extracted from the CNN outputs to the sixth Transformer layer, based on the observation that this layer contributes the most to the downstream speaker identification task (Chen et al., 2022).
@@
Conditional flow-matching (CFM) The input model in the CFM is composed of two feed forward networks (FFNs) and a linear layer, where each FFN has two linear layers with 512 hidden units and residual connection, with a ReLU activation and dropout rate of 0.05. Also, Layernorm is applied to the output of each FFN. The final linear layer projects the 512 dimensional feature to 256. The Transformer in the CFM has 8 layers and each layer has 8 heads with 64 dimensions, and 512 for the encoding dimension. We use Rotary positional embeddings (Su et al., 2024). The final output is projected to the 14 dimensional flow in articulatory feature space.
A.1.4 Training Details
@@
We train Sylber in two stages. The first stage is training with segment boundaries inferred from SDHuBERT which are extracted once at the beginning and fixed while in this stage. The second stage utilizes online segmentation using the teacher model’s outputs, by the algorithm in §3.2. Note that this training is only possible since our model exhibits features clean enough for our greedy segmentation to work. In the second stage, the L2 norm threshold is updated online by aggregating the statistics of speech and non-speech segments, and the merge threshold is randomly sampled from $[0.8,0.9]$. After the training, the norm threshold is fixed at 3.09 and the merge threshold is fixed at 0.8. See Appendix A.1.5 for details about the thresholding. Sylber is trained for 1.15M steps in the first stage and further trained for 500k steps in the second stage. We use a batch size of 64 and each data point is randomly cropped to be 5 seconds, following Cho et al. (2024b). The learning rate is set as 1e-4 with initial 500 warmup updates for the first stage and 5e-5 for the second stage. The EMA decay rates are set as 0.9995 and 0.9999 for the first and second stages, respectively. The second stage training improves performance in syllable detection and discovery (Appendix A.1.6).
For the CFM, the learning rate is fixed as 1e-4, with a batch size of 64 and 200k updates. For Articulatory Encodec and unit LM, we largely follow Cho et al. (2024c) and Lakhotia et al. (2021), respectively.
A.1.5 Thresholds Setting
@@
Thresholds in SDHuBERT segmentation For segmentation on SDHuBERT features, we apply the minimum cut algorithm introduced by Peng et al. (2023) and modified by Cho et al. (2024b). Following Cho et al. (2024b), the initial mask is obtained by thresholding norms of features from the eleventh layer of Transformer, where we normalize norms to be in $[0,1]$ and use 0.1 as threshold. The minimum cut refines each masked chunk to make it syllabic. Specifically, the algorithm conducts within segment agglomerative clustering witha preset number of clusters. This preset number is estimated by a pre-defined speaking rate. As this preset number of syllables may be larger than the number of segments, a post-hoc merging process merges adjacent segments with cosine similarity higher than a threshold, which we call the merge threshold. For tokenization experiments, we use a more sensitive segmentation configuration than the original setting to prevent loss of speech contents due to overly broad segments. Specifically, we halve the estimated syllable duration from 200ms to 100ms to cover speech with fast speaking rate, and increase the merge threshold from 0.3 to 0.4. However, the SDHuBERT is still sensitive to non-speech noise events. Therefore, we filter out segments with average absolute amplitude of waveform lower than 0.05
@@
Thresholds in Sylber greedy segmentation algorithm Unlike the thresholds in SDHuBERT, which are heuristically driven, we try to set the thresholds in our algorithm more principled way, especially in the second stage of training where the target segments are dynamically generated. We first set the norm threshold to be optimal boundary between signal (speech) and noise (non-speech), where the likelihoods of being signal and noise are equal. We assume both signal and noise distributions to be Gaussian and solve the equality condition. After the first training stage, we use the pseudo ground truth segments used for training to get the distribution of segment norms and non-segment norms in the dev split of LibriSpeech. To make the distribution reflects noise, we apply the noise augmentation as described in the denoising objective (Sec. A.1.2) to each sample. In the second training stage, we update the mean and variance of noise distribution using the non-segment portions of student outputs using an exponential moving average with a decay rate of 0.9999, while keeping the signal distribution the same as initially set. This results in the threshold of 3.09 after training. While these segments may not require such frequent updates in threshold, we implement this to try to keep it principled and empirically driven.
@@
On the other hand, we still remain largely heuristically driven in terms of setting our merge threshold. We use a particularly high threshold of 0.8 compared to 0.3 in the previous works. Such a high threshold for merging is effective in Sylber since the features are much cleaner than SDHuBERT. Instead setting this threshold to a fixed number, we sample a value from $[0.8,0.9]$ during the second stage training, which is somewhat arbitrarily set after visually inspecting multiple samples. We found that 0.7 also generally works fine but we select 0.8 as the threshold for the inference since that is the lowest number in the range we impose during training. In fact, the phoneme recognition experiment empirically proves that 0.8 is optimal when thresholds of 0.1 increments are tested (Table 8).
A.1.6 Effect of The Second Stage Training with Online Segmentation
To check the effectiveness of the second stage training with the online segmentation, we compare syllable detection and discovery metrics between the stage 1 and stage 2 models. As shown in Table 6, we observe some gain after the second stage training, especially in precision of the segmentation.
Model Syllable Detection Syllable Discovery
[…]
Sylber-Stage-1 73.7 69.2 71.4 75.6 63.2 43.9 5.24
Sylber-Stage-2 76.6 68.3 72.2 75.9 64.0 43.9 5.28
A.1.7 Effect of Denoising Objective
As demonstrated in left two panels in Figure 4, the syllabic structures are already highly visible without the denoising objective, indicating that the major learning source is self-segmentation distillation than the denoising objective. However, adding the denoising objective significantly improves robustness; otherwise, the model becomes highly sensitive to noisy audio as shown in the right two panels in Figure 4.
Figure 4: Frame-wise similarity matrix from with and without denosing objectives, using clean signal (left two panels) and noisy signal (right two panel). The orange waveform depicts the source noise we add to the clean speech signal.
A.2 Coding Efficiency with Duration-informed Tokenization
When we measure coding efficiency in §2, we ignore the duration information. Here, we recalculate the metrics by adding duration as separate token tagged to each speech token. Note that duration is counted as the number of frames, so it already lies on discrete space. We find that 99% of HuBERT tokens have duration less than 8, 7, and 6 with the vocab size of 50, 100, and 200, respectively. This means that the duration of each token can be coded by 3 bits. However, when BPE is applied, these 3 bits will be multiplied by the maximum number of units in subwords to count per-token duration bits, which is 10 to 16 depending on the vocab size and cluster granularity.
The syllabic tokens do not densely cover the frames. Therefore, the duration of subsequent silence can be tagged along with the duration of the tokens. 98% of syllabic tokens have duration less than or equal to 16 (4 bits). We can also keep the subsequent silence duration up-to 7 frames (3 bits) efficiently, and silence longer than 7 frames can be regarded as a separate “silence token”, adding one more token to the k-means codebook.
Taking all these into consideration, we measure the coding efficiency metrics with the duration-informed tokens as Table 7. Compared to Table 3, the gap between HuBERT-BPE and ours gets even larger, where we achieve around or more than 4×4$\times$4 × gains compared to HuBERT baselines. Moreover, even after appending duration tokens, we achieve significantly lower bitrates which are below or around 100.
Table 7: Coding efficiency of duration-informed tokens.
[…]
Vocab size Vocab size Vocab size
5K 10K 20K 5K 10K 20K 5K 10K 20K
HB50-BPE 7.45 6.82 6.30 449.26 418.24 392.36 0.0058 0.0062 0.0066
SDHB 5.84 112.73 118.58 124.42 0.0239 0.0228 0.0219
Sylber 4.76 91.80 96.56 101.32 0.0297 0.0284 0.0271
A.3 General Representational Power of Sylber
Though the universal utility of our model is not of our focus, we evaluate and benchmark downstream tasks using SUPERB (Yang et al., 2021). First of all, to find the optimal merge threshold, we train a phoneme recognition (PR) model with syllabic embeddings, where the merge threshold is sampled from $[0.3,0.9]$. The regular CTC based approach is not applicable to syllabic granularity, since it requires that the input length must be no shorter than the target length. Instead, we adopt RNN-T (Graves, 2012) which has no restriction on sequence length. To keep the model size similar to the PR model in SUPERB, we use a very simple, non-RNN transcriber, which is a Layernorm followed by two linear layers where the GELU activation function is applied to the first linear layer’s output. The output size of the first layer is set as 768 and set as the vocab size of phonemes, 73, for the second layer. The predictor network has a 3 layer LSTM with a hidden size of 1024, 0.1 dropout rate and Layernorm applied. The model is trained with the RNN-T implementation in PyTorch, and we use beam size of 5 for decoding. The learning rate is set as 0.001 and AdamW is used. The model is trained until no improvement is found in validation loss. We use LibriSpeech clean subsets (train-clean, dev-clean, and test-clean), which is the dataset used in SUPERB PR task setting. As results in Table 8, the merge threshold of 0.8 is selected and used throughout the SUPERB evaluation. This number coincides with the threshold we use in the main results as well. We use the code provided by S3PRL for the experiment. [^9] https://github.com/s3prl/s3prl
Table 8: Phoneme recognition on LibriSpeech (LS) dev-celan with different merge threshods.
Dataset PER $\downarrow$
Mthr=0.5 Mthr=0.6 Mthr=0.7 Mthr=0.8 Mthr=0.9
LS dev-clean 6.15 5.88 5.73 5.68 5.68
Sylber-Segment uses segment embedding after segmentation, with syllable granularity.
Sylber-Segment-Expand expands segment embedding to original length.
Table 9 compares these with a HuBERT base model, which has a comparable model size and trained on the same data. Since Sylber-Segment has a shorter sequence length than the target, thus making the CTC-based recognition task inapplicable, we replace the scores using the aforementioned RNN-T model, and we find a reasonable performance in PR as PER of 5.98, while ASR is lagging by large margin. As our model features are syllabic, this structure may need to be resolved to be converted to characters, adding additional layer of complexity than mapping phonemic features to characters which is hard to resolve in a limited resource setting.
Another notable point is that our models achieve higher keyword spotting accuracy (KS) and intent classification (IC) compared to the HuBERT base model in all 3 versions. This is aligned with the improved performance in language learning reported in §6.3. Also, there is a huge drop in speaker identity accuracy (SID) when our syllabic embedding is used, indicating that the speaker information is somewhat marginalized out.
Also, the failure in slot filling (SF) and automatic speech verification (ASV) by Sylber-Segment is attributed to the fact that S3PRL is tuned to lengthy input of speech representation with a regular sampling rate. Further investigation is required, for a proper application of syllabic embedding to those tasks.
Model PR KS IC SID ER ASR ASR (w/ LM) QbE SF ASV SD
[…]
Sylber-Segment 5.98 97.08 98.92 50.59 64.50 14.07 0.0139 13.21
A.4 Rhyming Word Pairs

### v2

@@ 7 Conclusion
We propose a novel self-supervised learning framework of speech, Sylber, that learns to transform speech waveform into a syllabic embedding that is well aligned with linguistic theories. Sylber offers promising potential for interpretable and efficient speech tokenization, and scalable and efficient spoken language modeling.
Limitations As we present our model more as a coding framework of speech, we largely put our focus on demonstrating efficiency and reconstruction quality. Therefore, our model is not yet suitable for universal speech representation, which the most speech SSL approaches aim for (Yang et al., 2021). We find that Sylber degrades in some SUPERB downstream tasks, which we believe is due to the parsimonious structure we are imposing. See Appendix A.2.4 and Table 12 for details and discussion. Also, the SUPERB protocol is optimally designed for frame-wise SSL; therefore, more investigation is needed on downstream architectures that better leverage the syllabic structure.
Acknowledgements
This research is supported by the following grants to PI Anumanchipalli — NSF award 2106928, BAIR Commons-Meta AI Research, the Rose Hills Innovator Program, UC Noyce Initiative at UC Berkeley, and Google Research Scholar, JP Morgan AI research Award. Special thanks to Shang-Wen (Daniel) Li and Abdelrahman Mohamed for valuable discussions and advice.
@@ A.1.1 Self-segmentation distillation
Given a speech audio signal $x$, we extract features, $M_{S}(x)=z^{S}$ and $M_{T}(x)=z^{T}$, where $M_{S}$ and $M_{T}$ are the student and teacher models, respectively. The unsupervised segmentation algorithm, Useg, outputs segment boundaries from $z$ as $\text{Useg}(z)=\{s\}^{N}$, where $N$ is the number of discovered segments, and $s\in\mathbb{N}^{2}$ denotes start and end frames of the segment, indexed as $s_{j,0}$ and $s_{j,1}$ for the $j$-th segment. We define an assignment function, $A(i)=j$, that gives the index of the segment, $j$, given a frame number, $i$, such that $s_{j,0}\leq i\leq s_{j,1}$. When there is no assignable segment, $A(i)=-1$, meaning $i$ is a non-speech frame. The segment-averaged feature, $v_{j}$, is defined by averaging across frames in the $j$-th segment: $v_{j}=\frac{1}{p-q}\sum_{k\in[p,q]}z_{k}$, where $(p,q)=(s_{j,0},s_{j,1})$. $v^{T}_{A(i)}$ is defined as the teacher’s segment-averaged feature of the segment that $i$-th frame belongs to, which is the target of the regression, which is zero for any non-speech frame: $v^{T}_{-1}=0$. Finally, the loss function of the proposed self-segmentation distillation is defined as $\mathcal{L}_{\text{SegDistill}}:=\sum_{i}||v^{T}_{A(i)}-z^{S}_{i}||_{2}^{2}$.
A.1.2 Greedy segmentation algorithm
@@
5:         if $\text{speech}_{i}$ and (no current segment or $\text{sim}(s_{i},\text{avg}(S_{k}))<M_{thr}$then
@@
8:              Add $s_{i}$ to current segment $S_{k}$ and update $\text{avg}(S_{k})$
@@
15:              if $\text{sim}(\text{avg}(S_{k}),\text{avg}(S_{k+1})\geq M_{thr}$ then
16:                  Merge $S_{k}$ and $S_{k+1}$
17:                  Continue
@@
19:              Define local search range from $a=\text{midpoint}(S_{k})$ to $b=\text{midpoint}(S_{k+1})$
20:              Get similarity to left or right segment $L=[\text{sim}(s_{i},\text{avg}(S_{k}))]_{a}^{b},R=[\text{sim}(s_{i},\text{avg}(S_{k+1}))]_{a}^{b}$
21:              Find optimal boundary $j^{*}$ = $\operatorname*{arg\,max}_{j}(\text{sum}(L_{1:j-1})+\text{sum}(R_{j:b-a}))$
22:              Update boundary to $j^{*}$
23:         end if
24:     end for
25:     return $S$
26:end procedure
The algorithm involves three linear passes through the audio embeddings (Algorithm 1). The first step thresholds all embeddings based on their L2 norm, distinguishing between speech and non-speech segments. Next, a monotonic agglomeration process iterates through the embeddings, grouping them into segments. At each time step, a frame is merged into the current segment if its cosine similarity with the segment’s average embedding exceeds a predefined threshold. This process runs in a single pass without constructing the entire similarity matrix by greedily starting a new segment whenever a frame falls below the threshold.
However, this greedy approach can introduce segmentation errors by slightly misaligning frame assignments. To correct this, a third pass refines segment boundaries. If adjacent segments have a cosine similarity above the merge threshold, they are merged. Otherwise, a local search range is defined from the midpoint of the previous segment to the midpoint of the following segment. Within this range, we compute the cosine similarity between each frame and the averages of the two segments. Instead of updating the averages dynamically, we use the original segment embeddings to keep the computational complexity linear. Finally, we determine the optimal boundary by maximizing the sum of cosine similarities between each frame and its assigned segment across all frames within the search range.
The time complexity is determined by the asymptotic number of dot product operations (or similarity computations), as these are the most computationally expensive steps in the segmentation algorithm.
A.1.3 Noise Augmentation
@@
For the denoising objective, we mix the input with a randomly sampled environmental sound or other speech audio. For mixing with environmental sound, we randomly select a clip from Reddy et al. (2021) and sample a 5 second clip from it. We first z-score the waveform and multiply by a factor sampled from $[0.05,0.7]$, and mix with the original speech audio. Note that the original speech is also z-scored. For mixing with other speech, we randomly select another clip in the batch and shift from left or right with a percentage sampled from $[0.4,0.7]$, to make sure the original speech holds the dominant information context in the mixture. The magnitude is also modulated by multiplying by a factor sampled from $[0.0,0.2]$. We apply this augmentation to 20% of the samples in the batch, and only to the inputs fed to the student model. Within the 20%, we have the source of noise be 75% environmental noise and 25% other speech.
A.1.4 Token-to-speech
@@
SPARC The target of our token-to-speech model is processed by SPARC (Cho et al., 2024c), which is composed of articulatory encoding and decoding. The encoding pipeline outputs 14 articulatory features at 50 Hz, which are composed of the XY coordinates of 6 articulators (lower incisor; upper and lower lips; tongue tip, blade and dorsum;), loudness, and pitch. These are interpretable and grounded representations of speech that are fully informative of speech contents (Cho et al., 2024c). The decoder, or articulatory vocoder, is a Hifi-GAN (Kong et al., 2020) conditioned on a speaker embedding inferred from a separate speaker encoder. Cho et al. (2024c) shows that SPARC successfully decomposes speech contents and speaker identity, by normalizing pitch to remove speaker specific pitch level. We replicate the implementation from Cho et al. (2024c), except that we change the layer of WavLM in the speaker encoder from the CNN outputs to the sixth Transformer layer, based on the observation that this layer contributes the most to the downstream speaker identification task (Chen et al., 2022). Following Kharitonov et al. (2021), the pitch is normalized by dividing it by the mean pitch and then taking its logarithm for the target of the token-to-speech model. The original pitch range is restored from the predicted normalized pitch by reversing the normalization process. For training SPARC, we use LibriTTS-R using a single A5000-24GB GPU.
@@
Conditional flow-matching (CFM) The input model in the CFM is composed of two feed forward networks (FFNs) and a linear layer, where each FFN has two linear layers with 512 hidden units and residual connection, with a ReLU activation and dropout rate of 0.05. Layernorm is applied to the output of each FFN. The final linear layer projects the 512 dimensional feature to 256. The Transformer in the CFM has 8 layers and each layer has 8 heads with 64 dimensions, and 512 for the encoding dimension. We use Rotary positional embeddings (Su et al., 2024). The final output is projected to the 14 dimensional flow in articulatory feature space. For training, the learning rate is fixed as 1e-4, with a batch size of 64 and 200k updates using a single A6000-48GB GPU.
Inference The SSL token embeddings are restored by the k-means codebooks, and expanded to the durations of the original segments. The non-speech frames are filled with zeros. For the case without quantization, the segment-averaged features are used. We use the extracted speaker embedding and mean pitch from the original speaker to synthesize speech from articulatory features predicted from tokens. We use the original segment durations for each token without predicting them since the duration information can be easily tokenized. (For example, duration can be tagged for each token. See Appendix A.2.3.) We remove randomness in the CFM to yield consistent generations for the evaluation purposes.
A.1.5 Sylber Training Details
@@
We train Sylber in two stages. The first stage is training with segment boundaries inferred from SDHuBERT which are extracted once at the beginning and fixed while in this stage. The second stage utilizes online segmentation using the teacher model’s outputs, by the algorithm in Section 3.2. Note that this training is only possible since our model exhibits features clean enough for our greedy segmentation to work. In the second stage, the L2 norm threshold is updated online by aggregating the statistics of speech and non-speech segments, and the merge threshold is randomly sampled from $[0.8,0.9]$. After the training, the norm threshold is fixed at 3.09 and the merge threshold is fixed at 0.8. See Appendix A.1.7 for details about the thresholding. The first-stage teacher is set and fixed as the initial student model, which is SDHuBERT with nine encoder layers. The second-stage teacher is then updated and fixed using the student model trained in the first stage. The model is trained for 115K steps in the first stage and further trained for 50K steps in the second stage. We use a batch size of 64 and each data point is randomly cropped to be 5 seconds, following Cho et al. (2024b). We show in Appendix A.2.9 that this does not significantly impact resynthesis for longer segments. The learning rate is set as 1e-4 with initial 500 warmup updates for the first stage and 5e-5 for the second stage. The second stage training improves performance in syllable detection and discovery (Appendix A.2.1). We used a single A6000-48GB GPU for training.
A.1.6 ULM Training Details
For training uLMs, we largely follow Hassid et al. (2024). When training on LibriLight, we use 96% of the data for training and 2% each is held out for validation and test. For each speech clip, we sample 100 tokens for training, which is roughly corresponding to 20 seconds. We use a single A6000-48GB GPU for training uLMs on LibriSpeech and two of them for training on LibriLight.
A.1.7 Thresholds Setting
@@
Thresholds in SDHuBERT segmentation For segmenting SDHuBERT features, we apply the minimum cut algorithm (MinCut) introduced by Peng et al. (2023) and modified by Cho et al. (2024b). Following Cho et al. (2024b), the initial mask is obtained by thresholding norms of features from the eleventh layer of the Transformer, where we normalize norms to be in $[0,1]$ and use 0.1 as threshold. The minimum cut refines each masked chunk to make it syllabic. Specifically, the algorithm conducts intra-segment agglomerative clustering with a preset number of clusters. This preset number is estimated by a pre-defined speaking rate. As this preset number of syllables may be larger than the number of segments, a post-hoc merging process merges adjacent segments with cosine similarity higher than a threshold, which we call the merge threshold. We use a more sensitive segmentation configuration than the original setting by Cho et al. (2024b) to prevent loss of speech contents due to overly broad segments. Specifically, we halve the estimated syllable duration from 200ms to 100ms to cover speech with fast speaking rate, and increase the merge threshold from 0.3 to 0.4. Note that the scores of SDHuBERT in Table 1 are those reported by Cho et al. (2024b), which use the original setting.
Non-speech Frames The non-speech frames are initially defined as“knocked out” frames by norm thresholding with SDHuBERT. However, we find that SDHuBERT is still sensitive to non-speech noise events. Therefore, we mark segments where the average absolute amplitude of z-scored waveform is lower than 0.1 as non-speech frames as well.
@@
Thresholds in Sylber greedy segmentation algorithm Unlike the thresholds in SDHuBERT, which are heuristically driven, we aim to set the thresholds in our algorithm in a more principled way, especially in the second stage of training where the target segments are dynamically generated. We first set the norm threshold to be optimal boundary between signal (speech) and noise (non-speech), where the likelihoods of being signal and noise are equal. Specifically, we assume both signal and noise distributions to be Gaussian and solve the equality condition. After the first training stage, we use the pseudo-ground truth segments used for training to get the distribution of segment norms and norms of non-speech frames in the dev split of LibriSpeech. To make the distribution reflect noise, we apply the noise augmentation as described in the denoising objective (Section A.1.3) to each sample. In the second training stage, we update the mean and variance of noise distribution using the non-segment portions of student outputs using an exponential moving average with a decay rate of 0.9999, while keeping the signal distribution the same as initially set. This results in the threshold of 3.09 after training.
@@
On the other hand, we still remain largely heuristically driven in terms of setting our merge threshold. We use a particularly high threshold of 0.8 compared to 0.3 in the previous works. Such a high threshold for merging is effective in Sylber since the features are much cleaner than SDHuBERT. Instead setting this threshold to a fixed number, we sample a value from $[0.8,0.9]$ during the second stage training, which is somewhat arbitrarily set after visually inspecting multiple samples. For inference, we select 0.8 as the threshold since it is the lowest number in the range we impose during training. While other threshold values (e.g., 0.7 or 0.9) generally work well, the phoneme recognition experiment in Appendix A.2.4 empirically proves that 0.8 is optimal when thresholds of 0.1 increments are tested (Table 11).
A.1.8 Syllable Segmentation Evaluation Details
The evaluation protocol for syllable detection and discovery follows the previous studies (Cho et al., 2024b; Komatsu & Shinozaki, 2024). We use the syllable labels of dev and test splits of LibriSpeech created by forced aligned phonemes and syllabication of them. Similar to SDHuBERT, Sylber interleaves silence frames between syllables. Thus, we use the front boundaries of the detected segments as the boundary predictions. Since this may incur some shifts in the detected boundaries, we use the dev split for finding the optimal constant shift to correct the boundaries, following the previous studies (Peng et al., 2023; Cho et al., 2024b; Komatsu & Shinozaki, 2024). The boundary prediction is considered a hit if it falls within a 50 ms tolerance window of the ground truth boundary. The scores are measured by precision (Pr), recall (Re), F1, and an additional R-value (R) that is introduced by Räsänen et al. (2009) to balance hit-rate and oversegmentation in measuring the quality.
For syllable discovery, the segment averaged embeddings are first clustered with 16384 centroids using k-means clustering and then the centroids are merged to 4096 using agglomerative (or hierarchical) clustering. We use the same setting as the previous studies (Peng et al., 2023; Cho et al., 2024b; Komatsu & Shinozaki, 2024) to make a fair comparison, whereas we use a separate clustering setting in the main tokenization experiments. The detected segments are mapped to the ground truth syllables by maximizing the temporal overlap between paired segments and syllables. We then measure the purity terms—cluster purity (CP), which indicates how purely each segment cluster is mapped to a syllable, and syllable purity (SP), which measures the reverse—as well as mutual information (MI). We refer to Section 4.1 of Komatsu & Shinozaki (2024) for formal definitions of these terms.
A.1.9 Rhyming Word Pairs